Hexamethylenediamine (C6H16N2) is one of the starting materials for the production of nylon. It can be prepared from adipic acid (C6H10O4) by the following overall reaction:


C6H10O4(l) + 2 NH3(g) + 4 H2(g) → C6H16N2(l) + 4 H2O(l)


What is the percent yield if 160.0 g hexamethylenediamine is made from 386.9 g adipic acid?


1
Expert's answer
2021-05-27T07:36:28-0400

Solution:

Balanced chemical equation:

C6H10O4(l) + 2NH3(g) + 4H2(g) → C6H16N2(l) + 4H2O(l)


The molar mass of adipic acid (C6H10O4) is 146.14 g/mol.

Therefore,

(386.9 g C6H10O4) × (1 mol C6H10O4 / 146.14 g C6H10O4) = 2.6475 mol


According to the equation above: n(C6H10O4) = n(C6H16N2)

n(C6H16N2) = n(C6H10O4) = 2.6475 mol


The molar mass of hexamethylenediamine (C6H16N2) is 116.21 g/mol.

Therefore,

(2.6475 mol C6H16N2) × (116.21 g C6H16N2 / 1 mol C6H16N2) = 307.66 g C6H16N2 = 307.7 g C6H16N2


307.7 g C6H16N2 - theoretical yield

160.0 g C6H16N2 - practical yield


%C6H16N2 = (160.0 g / 307.7 g) × 100% = 51.999% = 52.0%

%C6H16N2 = 52.0%


Answer: 52.0%